### v2

@@ 4.5.1 Settings
Figure 12: Distribution of scale values in MLICv2 on Kodak [37].
@@ 4.5.2 Analysis of Enhanced Transform
To demonstrate the superior capacity of our proposed transform, we compare the upper bounds of reconstruction quality achieved by the transforms in MLIC++[33], LIC-TCM[45], and our MLICv2. The results are shown in Fig. 10. Our MetaFormer-based design achieves better reconstruction quality than MLIC++ and performs comparably to LIC-TCM, highlighting its stronger representational power despite relying primarily on convolutional operations.
To demonstrate the effectiveness of the proposed simple token mixing block, we compare the transform module with the transform modules of Cheng’20 [16], SwinT-Charm [87], ELIC [30], and LIC-TCM [45] on rate-distortion performance. The entropy model is aligned with spatial autoregressive entropy model [50]. The results are presented in Table. 2. Our proposed transform achieves better performance than LIC-TCM while with about half forward MACs. When equiped with proposed transform, the model performance gets further enhanced. Since our baseline performance is very strong, the boost from the transform is slightly small relative to Table. 2.
@@ 4.5.3 Analysis of Enhanced Entropy Modeling
As shown in Table 4, disabling the selective compression module results in a 0.17% performance drop, which is acceptable, as the skipped elements—primarily zeros—typically have very low entropy despite accounting for the majority of all latent elements. By omitting the encoding of these zero elements, we can significantly accelerate both arithmetic encoding and decoding. To demonstrate the effectiveness of proposed guided selective compression on acceleration, the skip ratio, encoding acceleration and decoding acceleration on Kodak [37] and CLIC Pro Val [65] are presented in Table 6. As image quality improves, the proportion of zero elements that can be skipped decreases, which in turn reduces the acceleration gains for both encoding and decoding. However, high-resolution images usually contain more redundancy, allowing a larger number of elements to be skipped. This leads to greater potential for acceleration in such cases. In addition, we found that setting $\xi$ to 0.2, 0.3, or 0.4 has little effect on overall performance. This is mainly due to two reasons. First, as illustrated in Figure 12, most scale values fall within the range [0.1, 0.2], with only a small fraction between 0.2 and 0.4. Thus, using $\xi$ within this range is already effective in skipping most zero elements with low prediction error. The selective compression module only needs to correct a few wrongly skipped important elements. Second, we apply a weighted binary classification loss to train the selective compression module, using the ground-truth encoding mask as supervision.
@@
Xie’21 [75]
ELIC [30]
WACNN [88]
STF [88]
LIC-TCM [45]
FTIC [40]
MLIC+ [34]
MLIC++ [33]
WeConvene [23]
LALIC [20]
DCAE [47]
@@
Figure 13: Changes in loss, PSNR and bpp during the refinement of latent representation and side information of Kodim01 from Kodak dataset [37]. “Initial” denotes the original loss, psnr, bpp when using the latent representation and side information outputed by MLICv2.
@@
To validate the effectiveness of the proposed 2D RoPE, the 2D RoPE is replaced by original additive relative position embedding [46]. The performance degradation caused by the replacement demonstrates the superiority of the proposed 2D RoPE. Inspired by some 2D positional encoding designs [13, 57, 72], which split the feature tensor along the channel dimension and apply positional encoding along the horizontal axis to one half and the vertical axis to the other, we also experimented with a similar strategy. Specifically, we divided the input tensor along the channel dimension into two parts and applied RoPE-based rotations along the horizontal axis and vertical axis, respectively. We found that this alternative design does not lead to a performance difference. We also conduct experiments on the baseline model using 2D RoPE without learnable $\theta_{x}$ and $\theta_{y}$. This variant results in approximately a 0.4% performance drop.
@@ 4.5.4 Analysis of Iterative Latent Refinement
In case 1, the latent refinement is removed, which leads to around $4.89\%$ bit-rate increase compared to MLICv2+. The increased bit-rate demonstrates the effectiveness of the latent refinement for exploring performance potential. The changes of loss, PSNR and bpp are illustrated in Fig. 13. During the refinement, the loss and bpp generally first rise and then fall to lower than the initial loss and bpp, while the PSNR generally keeps rising.
@@ 5 Conclusion
In this paper, we present MLICv2 and MLICv2+, enhanced versions of the previous MLIC series that address observed limitations in transform design and entropy modeling. To overcome these limitations, we introduce efficient simple token mixing blocks that substantially improve representational capacity while maintaining computational efficiency. Our MetaFormer-based design achieves comparable performance to complex transformer alternatives with reduced overhead. For entropy modeling, we propose hyperprior-guided global context prediction, context reweighting mechanisms, and novel 2D Rotary Positional Embedding, collectively providing richer contextual information. The guided selective compression further accelerates encoding/decoding through intelligent zero-element skipping. Additionally, by employing SGA, we demonstrate the potential for performance improvements within our architectural framework. This validation analysis reveals the potential for further architectural improvements and provides valuable insights for future development directions. These innovations collectively establish new state-of-the-art performance benchmarks, with BD-Rate reductions of up to 20% compared to VTM-17.0 Intra while maintaining competitive computational complexity. Furthermore, to continue exploring the performance upper bounds via encoding-time scaling, we plan to investigate the transmission of update parameters [67, 59, 49] to the decoder side.
Limitations: Although our MLICv2 and MLICv2+ achieve state-of-the-art performance, several limitations warrant future attention. First, the performance potential exploration through SGA may lead to bitrate variations [84], requiring practical online encoder optimization strategies for precise bitrate control. Second, MLICv2/MLICv2+ may encounter cross-platform decoding issues [7, 29, 61, 79]. We plan to investigate model quantization techniques [28, 29, 61] and chain coding-based latent compression [79] to address deployment challenges.

### v3

@@ 4.5.1 Settings
Figure 13: Distribution of scale values in MLICv2 on Kodak kodak.
@@ 4.5.2 Analysis of Enhanced Transform
To demonstrate the superior capacity of our proposed transform, we compare the upper bounds of reconstruction quality achieved by the transforms in MLIC++jiang2023mlicpp, LIC-TCMliu2023learned, and our MLICv2. The results are shown in Fig. 11. Our MetaFormer-based design achieves better reconstruction quality than MLIC++ and performs comparably to LIC-TCM, highlighting its stronger representational power despite relying primarily on convolutional operations.
To demonstrate the effectiveness of the proposed simple token mixing block, we compare the transform module with the transform modules of Cheng’20 cheng2020learned, SwinT-Charm zhu2022transformerbased, ELIC he2022elic, and LIC-TCM liu2023learned on rate-distortion performance. The entropy model is aligned with spatial autoregressive entropy model minnen2018joint. The results are presented in Table. 2. Our proposed transform achieves better performance than LIC-TCM while with about half forward MACs. When equiped with proposed transform, the model performance gets further enhanced. Since our baseline performance is very strong, the boost from the transform is slightly small relative to Table. 2.
@@ 4.5.3 Analysis of Enhanced Entropy Modeling
As shown in Table 4, disabling the selective compression module results in a 0.17% performance drop, which is acceptable, as the skipped elements—primarily zeros—typically have very low entropy despite accounting for the majority of all latent elements. By omitting the encoding of these zero elements, we can significantly accelerate both arithmetic encoding and decoding. To demonstrate the effectiveness of proposed guided selective compression on acceleration, the skip ratio, encoding acceleration and decoding acceleration on Kodak kodak and CLIC Pro Val CLIC2020 are presented in Table 6. As image quality improves, the proportion of zero elements that can be skipped decreases, which in turn reduces the acceleration gains for both encoding and decoding. However, high-resolution images usually contain more redundancy, allowing a larger number of elements to be skipped. This leads to greater potential for acceleration in such cases. In addition, we found that setting $\xi$ to 0.2, 0.3, or 0.4 has little effect on overall performance. This is mainly due to two reasons. First, as illustrated in Figure 13, most scale values fall within the range [0.1, 0.2], with only a small fraction between 0.2 and 0.4. Thus, using $\xi$ within this range is already effective in skipping most zero elements with low prediction error. The selective compression module only needs to correct a few wrongly skipped important elements. Second, we apply a weighted binary classification loss to train the selective compression module, using the ground-truth encoding mask as supervision.
@@
Xie’21 xie2021enhanced
ELIC he2022elic
WACNN zou2022the
STF zou2022the
LIC-TCM liu2023learned
FTIC li2024frequencyaware
MLIC+ jiang2022mlic
MLIC++ jiang2023mlicpp
WeConvene fu2024weconvene
LALIC feng2025linear
DCAE lu2025learned
@@
Figure 14: Changes in loss, PSNR and bpp during the refinement of latent representation and side information of Kodim01 from Kodak dataset kodak. “Initial” denotes the original loss, psnr, bpp when using the latent representation and side information outputed by MLICv2.
@@
To validate the effectiveness of the proposed 2D RoPE, the 2D RoPE is replaced by original additive relative position embedding liu2021swin. The performance degradation caused by the replacement demonstrates the superiority of the proposed 2D RoPE. Inspired by some 2D positional encoding designs carion2020end, ramachandran2019stand, wang2020axial, which split the feature tensor along the channel dimension and apply positional encoding along the horizontal axis to one half and the vertical axis to the other, we also experimented with a similar strategy. Specifically, we divided the input tensor along the channel dimension into two parts and applied RoPE-based rotations along the horizontal axis and vertical axis, respectively. We found that this alternative design does not lead to a performance difference. We also conduct experiments on the baseline model using 2D RoPE without learnable $\theta_{x}$ and $\theta_{y}$. This variant results in approximately a 0.4% performance drop.
@@ 4.5.4 Analysis of Iterative Latent Refinement
In case 1, the latent refinement is removed, which leads to around $4.89\%$ bit-rate increase compared to MLICv2+. This significant performance improvement can be attributed to instance-specific adaptation. While our MLICv2 model is trained on a diverse dataset to achieve good average performance across various images, each individual image possesses unique characteristics such as texture complexity, edge distribution, and local redundancy patterns. ILR enables the latent representation $\hat{\boldsymbol{y}}$ and side information $\hat{\boldsymbol{z}}$ to be specifically optimized for each input image’s rate-distortion characteristics, effectively bridging the gap between generalized encoder mappings ($g_{a}$ and $h_{a}$) and input-specific optimal representations. The evolution of loss, PSNR, and bpp during refinement is illustrated in Fig. 14. During the initial refinement steps, both bit-rate and PSNR increase simultaneously, indicating that ILR adjusts latent values near quantization boundaries to better balance rate and distortion for the specific input. The subsequent phase exhibits decreasing bit-rate accompanied by further PSNR improvement, demonstrating that the refinement process successfully discovers a more optimal operating point in the rate-distortion space that remains inaccessible to the fixed encoder network. This validates the effectiveness of ILR in exploiting the full potential of our architectural framework.
@@ 5 Conclusion
In this paper, we present MLICv2 and MLICv2+, enhanced versions of the previous MLIC series that address observed limitations in transform design and entropy modeling. To overcome these limitations, we introduce efficient simple token mixing blocks that substantially improve representational capacity while maintaining computational efficiency. Our MetaFormer-based design achieves comparable performance to complex transformer alternatives with reduced overhead. For entropy modeling, we propose hyperprior-guided global context prediction, context reweighting mechanisms, and novel 2D Rotary Positional Embedding, collectively providing richer contextual information. The guided selective compression further accelerates encoding/decoding through intelligent zero-element skipping. Additionally, by employing SGA, we demonstrate the potential for performance improvements within our architectural framework. This validation analysis reveals the potential for further architectural improvements and provides valuable insights for future development directions. These innovations collectively establish new state-of-the-art performance benchmarks, with BD-Rate reductions of up to 20% compared to VTM-17.0 Intra while maintaining competitive computational complexity. Furthermore, to continue exploring the performance upper bounds via encoding-time scaling, we plan to investigate the transmission of update parameters tsubota2023universal, shen2023dec, lv2023dynamic to the decoder side.
Limitations: Although our MLICv2 and MLICv2+ achieve state-of-the-art performance, several limitations warrant future attention. First, the performance potential exploration through SGA may lead to bit-rate variations zhang2024practical, requiring practical online encoder optimization strategies for precise bit-rate control. Second, MLICv2/MLICv2+ may encounter cross-platform decoding issues balle2018integer, he2022post, shi2023rate, yang2025learned. We plan to investigate model quantization techniques han2015deep, he2022post, shi2023rate and chain coding-based latent compression yang2025learned to address deployment challenges.